Csc operations news feb 15 22
This presentation is the property of its rightful owner.
Sponsored Links
1 / 11

CSC Operations news Feb.15-22 PowerPoint PPT Presentation


  • 47 Views
  • Uploaded on
  • Presentation posted in: General

CSC Operations news Feb.15-22. M. Ignatenko (UCLA) CSC Operation and DPG meeting CERN February 22, 2012. CSC operations news. New TF firmware with 0-data suppression format is in action. No errors have been detected in many local and global runs.

Download Presentation

CSC Operations news Feb.15-22

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Csc operations news feb 15 22

CSC Operations newsFeb.15-22

M. Ignatenko (UCLA)

CSC Operation and DPG meeting

CERN

February 22, 2012


Csc operations news

CSC operations news

  • New TF firmware with 0-data suppression formatis in action.

    No errors have been detected in many local and

    global runs.

  • New ALCT firmware with 0-data suppression format.

    We have not seen much difference in the results

    relative to old firmware.


Dcc ddu

DCC+DDU

  • DCC and DDU firmware test during MWGR

    We hope that they reduce the rate of event with DCC

    and DDU out of sync (had been demonstrated in the high rate

    runs with random and radioactive triggers).

    At the same time the errors were detected even in the

    night Global run with L1ca rate 770 Hz.

  • A new DDU + DCC firmware are in FED Crate 1 and 4. Old firmware is in FED Crates 2 and 3.

    a. Increase timeout to 420 usec(wait time for DDU block).

    b. L1 mismatch counter doesn’t clear on resync (but clears on

    all other reset types).

    c. Gbit link to local DAQ (Spy channel) is functioning.

    d. Eliminate garbage packet on resync in spy channel.


Csc plan for mwgr

CSC Plan for MWGR

  • TFwants

    1. To demonstrate that we can run at 105 kHz

    without getting into busy state.

    2. To demonstrate a correct response on resync

    signal during busy state.

  • DCC+DDU

    - Global runs at different L1ac trigger rates

    (together with pixels; mostly random triggers)

    Long night cosmics run.

    What else?

    Automatic hard reset – can be also checked.


Dcc ddu test

DCC+DDU test

“High” rate run:

45 kHz mostly randoms

CSC+DT 180 Hz cosmics - bottom only.

ME-1/1/17, ME-1/1/19

And me+1/2/06 were ~100% in error.

1 OOS in the FED 756


Eprom firmware

EPROM firmware

  • The first ALCT lost EPROM information after 3 weeks since EPROM had been downloaded.

  • No CFEBs have been detected with EPROM problem so far (1 month of monitoring).


New hardware problems

New hardware problems

  • ME-1/1/19 bad ALCT data 100%

    Under investigation.

  • ME-1/1/17 bad ALCT data ~90%.

    Under investigation.

  • ME+1/2/06 bad ALCT data ~100%

    Under investigation.

  • ME+1/2/33 ALCT trigger bit (bits) is always on – hot ALCT.

    Under investigation. Wrong trigger stamp and timing.

    Backup solution – disable ALCT trigger and use CLCT only

    (the timing must be changed).

  • ME-3/1/03 unstable behavior – up to 63% ALCT data corrupted.

    Under investigation.

  • 2 TMBs had to be replaced in the VME-3/6 (lost VME communication)?

  • 1 DMB (me-1/121) lost VME communication. Fixed.


3 3 v tmb fuses

3.3 V TMB fuses

  • At least 2 TMBs have been found where 3.3 V fuses are loose in the fuse cradles.

    This can be another reason for “big”

    voltage drop on the 3.3 V lead.


Csc background

CSC background

Noisy HV segments

ME11 background reduced ~1.5 times after 1 month of HI run and 2 months without beam.

ME41 and ME42 increased due to beam pipe shield has been removed.


Access to the mini racks on the far side is fully blocked

Access to the mini racks on the far side is fully blocked


Shift information

Shift information

  • Next week (Feb. 29) we are starting regular DOC shifts.

  • Our DQM regular shifts are going to start one week later on March 07 (UCLA).


  • Login